Suppose the U.S. president wants to estimate the proportion of the
population that supports his current policy toward revisions in the
health care system. The president wants the estimate to be within 0.02
of the true proportion. Assume a 95% level of confidence. The
president’s political advisors found a similar survey from two years
ago that reported that 43% of people supported health care revisions.

A.How large of a sample is required? (Round intermediate values to 3
decimal points. Round up your answer to the next whole number.)

B.How large of a sample would be necessary if no estimate were
available for the proportion supporting current policy? (Round
intermediate values to 3 decimal points. Round up your answer to the
next whole number.)

please help
Answer #1

A. E = 0.02

p = 0.43

z = 1.96

n = (z/E)2*p*(1 - p)

n = (1.96/0.02)2*0.43*(1 - 0.43)

n = 2354

B. E = 0.02

p = 0.5

z = 1.96

n = (z/E)2*p*(1 - p)

n = (1.96/0.02)2*0.5*(1 - 0.5)

n = 2401

Please give me a thumbs-up if this helps you out. Thank you!